OpenStudy (anonymous):

for what value of x is g(x)=6?

OpenStudy (anonymous):

what is g(x)

OpenStudy (mayankdevnani):

\[x = x^{0}=1\]

OpenStudy (anonymous):

I'm not sure its in my Pre Cal homework its like an f(x) transformation

OpenStudy (mayankdevnani):

but you can't mentioned the x on R.H.S

OpenStudy (mayankdevnani):

@normag667

OpenStudy (anonymous):

ya but you have to define g(x) first like if \[g(x)=x ^{2}\] were the question you should do as follows g(x)=x ^2 so \[x ^{2}=6\] \[x=+/- \]

OpenStudy (anonymous):

\[\sqrt{6}\]

OpenStudy (mayankdevnani):

it can be -6

OpenStudy (mayankdevnani):